Please help me solve this substitution problem.
3(x-y)=15
2x+y=7
The question I have is does it matter which equation I use as the substitution? I used the 2nd equation and came up with a different answer than with the first.
First solution:
3(x-y)=15
2x+y=7 y=7-2x
3(x-7-2x)=15
3x-21-6x=15
-3x=36
x=-12
y=7-2x
y=7-2(-12)
y=31
Second solution:
3(x-y)=15 3x-3y=15
2x+y=7 3x=15+3y
x=5+y
2(5+y)+y=7
10+2y+y=7
3y=-3
y=-1
x=5+y
x=5+-1
x=4

which is correct and how do I know which equation to start with on these types of problems?

Answer by MathTherapy(10858)

3(x - y)= 15
2x + y = 7

2x + y = 7
y = 7 - 2x
Substituting this value for y in eq. (i) gives us:
3{x-(7-2x)} = 15
3(x – 7 + 2x) = 15
3x – 21 + 6x = 15
9x = 36
x = 4
Substitute this value for x in eq. (ii) 2x + y = 7. This give us:
2(4) + y = 7
y = - 1
You made an error when substituting the value of y (7 – 2x) into the 1st equation (see above). As seen, either equation can be substituted in order to find the values of x and y. So, the answer is (4, - 1).